# Zippo: Zipping Color and Transparency Distributions into a Single Diffusion Model

### 17 Mar 2024 v1

@@ Abstract
Beyond the superiority of the text-to-image diffusion model in generating high-quality images, recent studies have attempted to uncover its potential for adapting the learned semantic knowledge to visual perception tasks. In this work, instead of translating a generative diffusion model into a visual perception model, we explore to retain the generative ability with the perceptive adaptation. To accomplish this, we present Zippo, a unified framework for zipping the color and transparency distributions into a single diffusion model by expanding the diffusion latent into a joint representation of RGB images and alpha mattes. By alternatively selecting one modality as the condition and then applying the diffusion process to the counterpart modality, Zippo is capable of generating RGB images from alpha mattes and predicting transparency from input images. In addition to single-modality prediction, we propose a modality-aware noise reassignment strategy to further empower Zippo with jointly generating RGB images and its corresponding alpha mattes under the text guidance. Our experiments showcase Zippo’s ability of efficient text-conditioned transparent image generation and present plausible results of Matte-to-RGB and RGB-to-Matte translation.
@@
Figure 1: An illustration of comparison between our proposed Zippo and a separate text-to-image diffusion model and image matting model. (a) text-to-image diffusion model generates a RGB image acccording to the input text prompt; (b) image matting model predict alpha matte from the input RGB image; (c) Zippo actually models the joint distribution of RGB and transparency information, thus acts as a multi-task learner. Specifically, besides the capabilities of RGB-to-Matte and Matte-to-RGB, Zippo can generate a pair of RGB image and alpha matte, thus compositing a transparent image.
@@ 3.1.1 Datasets
Zippo is trained on two datasets AM2K [24] and VITON-HD [9]. Considering the important human-centric characteristics of generating transparent portrait images for vast applications (e.g., film scene/character composition, professional portrait layer composition, and layout design), it is essential to gather high-quality unobstructed portrait images for training Zippo. While P3M10K [22] is a commonly used dataset in portrait matting, we have opted not to utilize it due to the privacy masks presented on human faces, which could be harmful to the quality of portrait generation. Therefore, we use an open-source virtual try-on dataset VITON-HD, which has 10k high quality human-centered images without masks on the face. For high-quality alpha matte labels, we first choose trimap-free portrait matting model P3M [22] to obtain the coarse alpha labels. Then we predict segmentation masks leveraging the strong semantic prior of SAM trained on 11B data [20] and filter out those that are deviated from the alpha labels. In concert, we quantize the alignments between alpha matte (binarized by a threshold $h=25$) and segmentation masks through computing IoU metric, and we only collect alpha labels that have IoU result greater than the threshold $\hat{h}$=0.9 [8]. Finally, we estimate a trimap from the filtered alpha mask as an auxiliary input for trimap-based matting model [29] which outputs the final fine-grained alpha labels. To explore the generalization of our model, we also adapt AM2K as another training data which has 2k high-quality animal images with human-labeled high quality alpha mattes.
@@ 3.2 Main Results
In this section, we present the results of Matte-to-RGB, RGB-to-Matte and joint generation (transparent image generation) on the datasets of AM2K and VITON. In Figs. 4 and 5, we present the results of joint generation of RGB images and corresponding alpha mattes. From the results, we can conclude that Zippo generalizes well to the cases of both animal and human. It is noteworthy that Zippo can not only generate photo-realistic RGB images but also simultaneously output its corresponding fine-grained alpha matte. By combining these two outputs together, we show the composed transparent images to verify the accurate alignment between RGB and matte and the content adherence to the text prompt. For example, in Fig. 4, Zippo generated a furry dog with accurate reasonable details, even predicting alpha transparency value for each strand of dog fur.
Additionally, we provide the results of Matte-to-RGB and RGB-to-Matte on AM2K and VITON in Figs. 6, 7, 8, 9. In Figs. 6 and 7, we take the alpha masks as the structural inputs and use the text prompts to specify the contents in the images, e.g., the classes of animals in the images and the clothes worn by the people. As can be concluded, our Zippo is capable of synthesizing animals and humans (Matte-to-RGB) with structural adherence to the input alpha mattes and plausible photo-realistic contents with fancy details. In Figs. 8 and 9, we take the RGB images as input and use Zippo to predict the corresponding alpha mattes (RGB-to-Matte). As shown in the Fig. 8, Zippo is able to estimate the alpha mattes for input images of arbitrary resolutions and successfully mat the foreground objects with detailed masks.
Figure 6: Results of Matte-to-RGB generation on AM2K.
Figure 7: Results of Matte-to-RGB generation on VITON.
Figure 8: Results of RGB-to-Matte prediction on AM2K.
Figure 9: Results of RGB-to-Matte prediction on VITON.
@@ 4 Related Work
Text-to-Image Diffusion Models Text-to-Image diffusion models are a class of diffusion-based image generative models that aim to generate images with content semantically aligned with the given text prompt [13]. Text-to-image diffusion models start from random Gaussian noises and take text prompt as the input and gradually de-noise the noise under the guidance to output an image in high fidelity. Recently, many text-to-image diffusion models have been proposed by giant companies as their magical generation fidelity and promising future applications. Rombach et al. trained a latent diffusion model called Stable Diffusion [33] on a large-scale image and text dataset, LAION-5B, and demonstrated previously unachievable image synthesis quality. In contrast to Stable Diffusion, DeepFloyd IF operates directly in the pixel space.
Utilizing pre-trained generative prior for perceptive tasks. Generative models are trained on large-scale data and thus could be viewed as pre-trained models with strong prior. Thus, increasingly works begin to explore how to leverage the priors of generative models for perceptual tasks. Recent works [3, 10] demonstrate that generative models encode characteristics of the scene. By finding latent variable offsets, using LoRA [15], etc., generative models can directly produce intrinsic images like surface normals, depth, albedo, etc. LDMSeg [40] proposed an image-conditioned sampling process to make diffusion models directly output the panoptic segmentation results. UniGS [32] proposes location-aware color encoding and decoding strategies, allowing diffusion models to support referring segmentation and entity segmentation. Marigold [17] fine-tunes diffusion model on limited synthetic data, enabling it to support affine-invariant monocular depth estimation and exhibit strong generalization performance. However, Marigold is time-consuming due to the need for multiple iterations of denoising. Additionally, the Gaussian noise leads to inconsistent results across inferences, requiring aggregation over multiple inferences.
@@
Image matting. Image matting is a typical ill-posed problem to estimate the foreground’s alpha matte from a single image. From the perspective of auxiliary input, image matting can be categorized into two types of methods. Trimap-based image matting methods require an auxiliary input trimap to mitigate the uncertainty [47, 46, 29, 27]. Meanwhile, researchers attempt to learn an automatic end-to-end model without any auxiliary inputs namely, trimap-free image matting methods [22, 18, 49, 4]. Previous methods include affinity based methods [2, 21] and sampling-based methods [12, 42]. Although trimap-free image matting methods are more applicable and flexible, they still fall behind the trimap-based matting methods according to their performance. Therefore, we design our pesudo-labeling roadmap by first obtaining the coarse alpha mask from trimap-based method, then computing a trimap according to the alpha mask and input it to the trimap-based method for a fine-grained alpha matte as our final pesudo-labels. It is worth noting that, previous image matting methods are typical discriminative ones by regressing the estimation results, while we try to solve this problem in a generative manner, by formulating the image matting problem as learning a conditional alpha matte distribution given the input images.

### 19 Mar 2024 v2

@@ Abstract
Beyond the superiority of the text-to-image diffusion model in generating high-quality images, recent studies have attempted to uncover its potential for adapting the learned semantic knowledge to visual perception tasks. In this work, instead of irreversibly translating a generative diffusion model into a visual perception model, we explore to retain the generative ability with the perceptive adaptation. To accomplish this, we present Zippo, a unified framework for zipping the color and transparency distributions into a single diffusion model by expanding the diffusion latent into a joint representation of RGB images and alpha mattes. By alternatively selecting one modality as the condition and then applying the diffusion process to the counterpart modality, Zippo is capable of generating RGB images from alpha mattes and predicting transparency from input images. In addition to single-modality prediction, we propose a modality-aware noise reassignment strategy to further empower Zippo with jointly generating RGB images and its corresponding alpha mattes under the text guidance. Our experiments showcase Zippo’s ability of efficient text-conditioned transparent image generation and present plausible results of Matte-to-RGB and RGB-to-Matte translation.
@@
Figure 1: An illustration of comparison between our proposed Zippo and a separate text-to-image diffusion model and image matting model. (a) text-to-image diffusion model generates a RGB image acccording to the input text prompt; (b) image matting model predict alpha matte from the input RGB image; (c) Zippo actually models the joint distribution of RGB and transparency information, thus acts as a multi-task learner. Specifically, besides the capabilities of RGB-to-Matte and Matte-to-RGB, Zippo can generate a pair of RGB image and alpha matte simultaneously, thus compositing a transparent image.
@@ 3.1.1 Datasets
Zippo is trained on two datasets AM2K [22] and VITON-HD [9]. Considering the important human-centric characteristics of generating transparent portrait images for vast applications (e.g., film scene/character composition, professional portrait layer composition, and layout design), it is essential to gather high-quality unobstructed portrait images for training Zippo. While P3M10K [20] is a commonly used dataset in portrait matting, we have opted not to utilize it due to the privacy masks presented on human faces, which could be harmful to the quality of portrait generation. Therefore, we use an open-source virtual try-on dataset VITON-HD, which has 10k high quality human-centered images without masks on the face. For high-quality alpha matte labels, we first choose trimap-free portrait matting model P3M [20] to obtain the coarse alpha labels. Then we predict segmentation masks leveraging the strong semantic prior of SAM trained on 11B data [18] and filter out those that are deviated from the alpha labels. In concert, we quantize the alignments between alpha matte (binarized by a threshold $h=25$) and segmentation masks through computing IoU metric, and we only collect alpha labels that have IoU result greater than the threshold $\hat{h}$=0.9 [8]. Finally, we estimate a trimap from the filtered alpha mask as an auxiliary input for trimap-based matting model [27] which outputs the final fine-grained alpha labels. To explore the generalization of our model, we also adapt AM2K as another training data which has 2k high-quality animal images with human-labeled high-quality alpha mattes.
@@ 3.2 Main Results
In this section, we present the results of Matte-to-RGB, RGB-to-Matte and joint generation (transparent image generation) on the datasets of AM2K and VITON. In Figs. 5 and 4, we present the results of joint generation of RGB images and corresponding alpha mattes. From the results, we can conclude that Zippo generalizes well to the cases of both animal and human. It is noteworthy that Zippo can not only generate photo-realistic RGB images but also simultaneously output its corresponding fine-grained alpha matte. By combining these two outputs together, we show the composed transparent images to verify the accurate alignment between RGB and matte and the content adherence to the text prompt. For example, in Fig. 5, Zippo generated a furry dog with accurate reasonable details, even predicting alpha transparency value for each strand of dog fur.
Figure 6: Results of Matte-to-RGB generation on AM2K. The first and third columns are input alpha mattes and the second and the forth columns are the generated results by Zippo. The text prompts are "a dog", "a tiger", "a rabbit" and "a deer", respectively.
Figure 7: Results of Matte-to-RGB generation on VITON. The first and third columns are the input alpha matte, and the second and fourth columns present the images in different resolutions generated by Zippo. Zoom in for the detailed alignments. For the first column, the text prompts are "a man wearing a black hoodie", and "a woman wearing a blue hoodie". For the third column, the text prompts are "a man", "a woman" and "a woman wearing a yellow t-shirt."
Additionally, we provide the results of Matte-to-RGB and RGB-to-Matte on AM2K and VITON in Figs. 6, 7, 8, 9. In Figs. 6 and 7, we take the alpha masks as the structural inputs and use the text prompts to specify the contents in the images, e.g., the classes of animals in the images and the clothes worn by the people. As can be concluded, our Zippo is capable of synthesizing animals and humans (Matte-to-RGB) with structural adherence to the input alpha mattes and plausible photo-realistic contents with fancy details. In Figs. 8 and 9, we take the RGB images as input and use Zippo to predict the corresponding alpha mattes (RGB-to-Matte). As shown in Fig. 8, Zippo is able to estimate the alpha mattes for input images of arbitrary resolutions and successfully mat the foreground objects with detailed masks.
Figure 8: Results of RGB-to-Matte prediction on AM2K. The first column shows input images in the test set, the second column shows predicted alpha mattes, and the third column shows the ground truth. It’s worth noting that Zippo presents great details compared to the ground truth alpha matte.
Figure 9: Results of RGB-to-Matte prediction on VITON. The first and third column is the input image, and the second and fourth column is the predicted alpha matte.
@@ 4 Related Work
Text-to-Image Diffusion Models Text-to-Image diffusion models are a class of diffusion-based image generative models that aim to generate images with content semantically aligned with the given text prompt[12]. Text-to-image diffusion models start from random Gaussian noises taking text prompt as the input and gradually denoise the noise under the guidance to output an image in high fidelity. Recently, many text-to-image diffusion models have been proposed by giant companies as their magical generation fidelity and promising future applications. Rombach et al. trained a diffusion model in latent space called Stable Diffusion[31] on a large-scale image and text dataset, LAION-5B, and demonstrated previously unachievable image synthesis quality.
Generative models for visual perceptive tasks Generative models capture strong images prior by training on large-scale text-image datasets. Recently, increasingly works explored to leverage the priors of generative models for visual perceptual tasks. Recent works [14, 42, 24] demonstrate that generative models are capable of solving dense prediction tasks like depth estimation, segmentation, and normal estimation. By finding latent variable offsets[3], generative models can directly produce intrinsic images like surface normals, depth, albedo, etc. Marigold [15] fine-tunes unet backbone of diffusion models to perform affine-invariant monocular depth estimation and exhibit strong generalization capability. Meta prompt [39] proposes to harness generative models for visual perceptive tasks by introducing learnable embeddings to the pre-trained generative models to extract proper features for visual perceptive tasks.
@@
Image matting. Image matting is a typical ill-posed problem to estimate the foreground’s alpha matte from a single image. From the perspective of auxiliary input, image matting can be categorized into two types of methods. Trimap-based image matting methods require an auxiliary input trimap to mitigate the uncertainty[45, 44, 27, 25]. Meanwhile, researchers attempt to learn an automatic end-to-end model without any auxiliary inputs namely, trimap-free image matting methods[20, 16, 47, 4]. Previous methods include affinity based methods[2, 19] and sampling-based methods[11, 40]. Although trimap-free image matting methods are more applicable and flexible, they still fall behind the trimap-based matting methods according to their performance. Therefore, we design our pesudo-labeling roadmap by first obtaining the coarse alpha mask from trimap-based method, then computing a trimap according to the alpha mask and input it to the trimap-based method for a fine-grained alpha matte as our final pesudo-labels. It is worth noting that, previous image matting methods are typical distcriminative ones by regressing the estimation results, while we try to solve this problem in a generative manner, by formulating the image mattiing problem as learning a conditional alpha matte distribution given the input images. the input images.